OpenStudy (anonymous):

The length of two sides of a right triangle are given. Find the length of the third side. Round to the nearest length if necesary. Legs: 2in and 15in

OpenStudy (anonymous):

Use the Pythagorean Theorem to solve for the hypotenuse. 2^2 +15^2=c^2

OpenStudy (anonymous):

I got 15.1 but its not in the choices

OpenStudy (anonymous):

Do the directions say to round to the nearest "tenth?" Also, are you sure that both given sides are the legs?

OpenStudy (anonymous):

Yes it does say round to the nearest tenth and no I'm not sure

OpenStudy (anonymous):

yeah... is the side that is 15 a leg or the hypotenuse?

OpenStudy (anonymous):

if the 15 is a hypotenuse, then the equation will be 2^2 + c^2 = 15^2... solve for c here...